OpenStudy (anonymous):

52 ÷ (4 + 1) 1 3 4 5

OpenStudy (anonymous):

@musicdramalife00 just this last prob.

OpenStudy (rockinhood):

It depends if you use PEMDAS Because if you do, start with the parenthesis (4 + 1 = 5) then you divide 52 / 5 = 10.4

OpenStudy (anonymous):

but 10.4 isnt there

OpenStudy (anonymous):

are you sure thats the right problem cause its 52/5 wait is there a x somewhere?

OpenStudy (rockinhood):

I don't know if I can help you then.

OpenStudy (anonymous):

you can't distribute it would end up being 52/4 + 52/1 = 13+52 = 65 As your problem is written, the answer is 10.4

OpenStudy (anonymous):

what is the 1 2 3 4 5 ???

OpenStudy (anonymous):

can you like take a pic of the thing or something

OpenStudy (anonymous):

no the problem is actually 5^2 ÷ (4 + 1)

OpenStudy (anonymous):

=.=

OpenStudy (anonymous):

the 12345 are the answers to choose from @ehuman

OpenStudy (anonymous):

oh